OpenStudy (anonymous):

The ticket office at Orchestra Center estimates that if it charges x dollars for box seats for a concert, it will sell 50 - x box seats. The function S = 50x - x2 gives the estimated income from the sale of box seats. Graph the function, and use the graph to find the price for box seats that will give the greatest income

OpenStudy (anonymous):

Correct. S = x * (50 - x) = 50x - x^2 You may want to know what the maximum income S(x) is: dS/dx = 50 - 2x = 0; x = 25 S(25) = 25 * 25 = 625 The maximum income is $625. The ticket price for that is $25.

OpenStudy (anonymous):

for a.b.c.d they have a graph

OpenStudy (anonymous):

I'm not too sure on the graph but just look at my final answer and look for a graph that proves that answer.

OpenStudy (anonymous):

yes it did.

OpenStudy (anonymous):

go to desmos.com and input the function for the graph. it will give you a clear look at to what the graph should look like.

OpenStudy (anonymous):

then wat

OpenStudy (anonymous):

best responce plz